OpenStudy (anonymous):

I need to find out if this question is always, sometimes or never true> Can someone help me

OpenStudy (anonymous):

This is what I have so far

OpenStudy (anonymous):

6+8x-9=11x+14-3x (Equation) +3x +3x _______________________________________ 6+8x-9=14x+14 +9 +9 _______________________________________ 15+8x=14x+14 -14 -14 _______________________________________ 1+8x=14x -8x -8x _______________________________________ 1=4x So what do i do next?